# CGKN: A Deep Learning Framework for Modeling Complex Dynamical Systems and Efficient Data Assimilation

### v1

@@ 3 Numerical Experiments
    The true model. This corresponds to the governing equation of the complex dynamical systems in Eq. (LABEL:eq:DS_u1u2). State predictions are made by averaging ensemble simulations from any given initial state, while DA results are obtained using the ensemble Kalman-Bucy filter (EnKBF) [51] based on data from observed states. In this work, the EnKBF is employed exclusively by this model for DA, while other models use the analytic DA formulae from Eq. (LABEL:eq:CGKN_CGF) to ensure a fair comparison. The test results from the true model serve as a benchmark for optimal performance.
@@
    Simplified Koopman network (Simplified KoopNet). This model applies the standard Koopman theory to the dynamics of the unobserved state $\mathbf{u}_{2}$, resulting in a linear governing equation for the latent state variables $\mathbf{v}$, rather than the nonlinear but conditionally linear structure of $\mathbf{v}$ found in the CGKN model. It is worth highlighting that this model still retains a strongly nonlinear structure in the dynamics of $\mathbf{u}_{1}$. The detailed form of the simplified Koopman network is:
@@
    $\frac{\mathop{}\!\mathrm{d}\mathbf{v}}{\mathop{}\!\mathrm{d}t}=\mathbf{F}_{2}+\mathbf{G}_{2}\mathbf{v}+{\bm{\sigma}}_{2}\dot{\mathbf{W}}_{\mathbf{2}},$
    where $\mathbf{F}_{2}\in\mathbb{R}^{d_{\mathbf{v}}}$ and $\mathbf{G}_{2}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ are two parameter matrices with other components staying the same as the CGKN model in Eq. (2.3). This standard application of the Koopman operator remains to impose a conditional Gaussian structure on the model, but the linear governing equation for $\mathbf{v}$ overlooks the contribution of the observed state $\mathbf{u}_{1}$. Numerical results show that, in comparison to the simplified KoopNet model, the CGKN model’s conditional linear structure in the governing equation of $\mathbf{v}$ significantly enhances its expressiveness. The comparison between the CGKN and the simplified KoopNet model highlights the crucial role of the conditional linear, yet still nonlinear, structure in the governing equation of $\mathbf{v}$ in the former.
@@
    Conditional Gaussian regression (CG-Reg). This is a regression model from system identification that involves two steps: (i) identifying significant basis functions from a library of system dynamics using a causal-based metric called causation entropy [96, 97, 98, 99, 100, 101], and (ii) estimating the coefficients of these significant basis functions via the least squares method or maximum likelihood estimation. The library is specifically designed to ensure that the resulting regression model conforms to the structure of conditional Gaussian nonlinear system in Eq. (2.3), allowing the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) to still be applied for DA [21]. In this work, this specially designed library is referred to as the CG-library, which requires that the unobserved state variables $\mathbf{u}_{2}$ appear linearly in all the basis functions. The detailed form of the CG-Reg model is as follows:
@@
    where $\mathbf{u}$ is the original system state variables including both the observable and unobservable ones. This DNN model does not follow the conditional Gaussian structure, and therefore, the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) cannot be applied for efficient DA. Numerical results demonstrate that the DNN model achieves state forecast performance comparable to that of the CGKN model; however, the additional key advantage of the CGKN model lies in its unique efficiency in DA.
For all numerical test cases, we assume that the training data includes both observable and unobserved state variables, while the testing data contains only the observed state variables. To evaluate the performance of models in both state prediction and DA across all examples, the normalized root mean squared error (NRMSE) is used:
$\textrm{NRMSE}:=\frac{1}{M}\sum_{i=1}^{M}\frac{\sqrt{\frac{1}{N}\sum_{n=1}^{N}\big{(}\mathbf{x}_{i}^{\dagger}(t_{n})-\mathbf{x}_{i}(t_{n})\big{)}^{2}}}{\textrm{std}(\mathbf{x}_{i}^{\dagger})},$ (3.4)
where $\mathbf{x}^{\dagger}(t_{n})\in\mathbb{R}^{M}$ is the true variables at the time step $t_{n}$, $\mathbf{x}(t_{n})$ is the approximated variables, $\mathbf{x}_{i}(t_{n})$ denotes the $i$-th variable, and $\textrm{std}(\cdot)$ corresponds to the standard deviation estimated by the samples from different time steps $t_{n}=1,...,N$. The NRMSE between true state variables and predictive onces is referred to as the forecast error, and the NRMSE between the true state variables and estimated state variables from DA is referred to as the DA error. Both forecast error and DA error are calculated using test data. The forecast error is calculated based on the same total forecast time steps in training. In contrast, the DA error is calculated over the whole time range of the test dataset. The test results for all models and all numerical examples are summarized in Table 3.1.
Table 3.1: Test results of state forecast and DA of all models for each example. The errors are the normalized root mean squared error (NRMSE) between true values and approximated values.
Models Examples PSBSE Lorenz 96 ENSO PDE Forecast Error DA Error Forecast Error DA Error Forecast Error DA Error True Model 2.4566e-01 6.9466e-01 4.3874e-02 6.9582e-02 6.5406e-01 CGKN 2.8389e-01 7.2776e-01 6.3738e-02 9.4883e-02 6.4762e-01 1.3518e-01 Simplified KoopNet 3.7600e-01 7.5069e-01 3.0589e-01 2.7252e-01 6.8473e-01 2.3490e-01 CG-Reg 2.7766e-01 1.2884e+00 2.7631e-01 3.5670e-01 9.8751e-01 NaN DNN 2.4728e-01 4.5174e-02 6.4845e-01
@@
The simulation results of 1000 time units are generated and provide the data with a sub-sampling step size $\Delta t=0.01$ time unit, where the first 800 units are utilized for training and the remaining 200 units are employed for testing. Figure 3.1 shows part of the simulation of PSBSE in panel (a), and the probability density functions (PDFs) and auto-correlation functions (ACFs) in panels (b) and (c), respectively. It can be seen in Fig. 3.1 that the true system demonstrates several complex features such as intermittency, extreme events, and non-Gaussian statistics. To compare the DA capability of each model, $x$ is set as the observed state, while $y$ and $z$ are set as unobserved state in this test case.
@@
where the latent state variables $\mathbf{v}$ are transformed from unobserved state variables $[y,z]^{\mathtt{T}}$ via the encoder $\bm{\varphi}$. The dimension of $\mathbf{v}$ is set as 10 in this test case, and the sub-networks in the CGKN model are $\bm{\varphi}:\mathbb{R}^{2}\mapsto\mathbb{R}^{10}$, $\bm{\psi}:\mathbb{R}^{10}\mapsto\mathbb{R}^{2}$, and $\bm{\eta}:\mathbb{R}^{1}\mapsto\mathbb{R}^{121}$ with number of parameters $\bm{\theta_{\bm{\varphi}}}$, $\bm{\theta_{\bm{\psi}}}$, and $\bm{\theta_{\bm{\eta}}}$ be 8778, 8770, and 5785 respectively. The training settings of the CGKN include: state prediction time range $[t_{0},t_{N_{s}}]$ is $0.1$ time units, and the DA time range $[t_{0},t_{N_{l}}]$ is $50$ time units with the warm-up period set as $t_{N_{b}}=3$ time units. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=\lambda_{\mathbf{v}}=1/d_{\mathbf{u}}$, and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is set as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler.
@@
We investigate the performance of both state prediction and DA of all models on the test data and report the results in Table 3.1 (under the column PSBSE). The forecast error is calculated by 0.1 time units as the forecast time range and the DA Error is based on 200 time units as the DA time range. For the true model, the ensemble number is set to 100 for both the ensemble mean forecast and the EnKBF. The error in the true model primarily arises from intrinsic random noise in the system, along with sampling errors of the ensemble methods. Consequently, the test results of the true model serve as a benchmark for evaluating the performance of other models. Among all other models, the CGKN model provides the best overall performance in both state forecast and DA. Although the DNN model slightly outperforms the CGKN model for state forecast, its black-box nonlinear architecture demands more computationally expensive techniques for DA, while the CGKN model can exploit the analytic formulae of Eq. (LABEL:eq:CGKN_CGF) for efficient DA. The forecast error from the CG-Reg model is slightly better than that of the CGKN model. The main reason is that, except for the nonlinear $yz$ term, all other terms in the true system of Eq. (3.5) can still be successfully identified by the CG-Reg model, if all those terms have been included in the CG-library. On the other hand, the CG-Reg model provides the worst performance of DA, indicating that the nonlinear term of $yz$ in the governing equation of $x$ plays an important role in the probabilistic relation between the observed state variable $x$ and the unobserved ones $y$ and $z$. Regarding the simplified KoopNet model, it is expected that the state forecast error will be noticeably larger than that of the CGKN model, primarily due to the omission of the contribution from the observed state variable $x$ in the governing equations of the unobserved state variables.
@@
Figure 3.2 presents the DA results from the four models for estimating the unobserved states $[y,z]^{\mathtt{T}}$ based on the observed state $x$. To obtain the posterior mean estimates, the EnKBF is employed for the true model, while the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) are used for the other three models. The uncertainty regions for the true model and the CG-Reg model are derived from the EnKBF and the covariance evolution formula in Eq. (LABEL:eq:CGKN_CGF), respectively. In contrast, the uncertainty regions for the CGKN and the simplified Koopman network are obtained using the residual analysis method described in Section 2.1.2. The comparable performance of the posterior mean estimation to that of the true model underscores the effectiveness of the CGKN in DA. Furthermore, the estimated uncertainty region encompasses most of the results from the true model, even though it remains narrower than that produced by the true model with EnKBF. In this numerical test case, the DA results from the simplified KoopNet model shown in Fig. 3.2(c) are comparable to those of the CGKN model, while the results from the CG-Reg model in Fig. 3.2(d) are noticeably inferior to those of the other models, particularly for the unobserved state variable $z$.
@@
Figure 3.2: DA results of the projected stochastic Burgers–Sivashinsky equation. EnKBF is used for the true model and analytic formulae are used for the other three models. The uncertainties are indicated by the grey-colored regions, which correspond to two estimated standard deviations from the posterior mean. The uncertainty area of the CGKN model and the simplified KoopNet model are estimated based on the method of residual analysis.
@@ 3.2 Lorenz 96 system: A forty-dimensional chaotic system
where $I=40$, $F=8$ and $\sigma=0.5$. With chaotic and wave-like behavior, the Lorenz 96 system can be regarded as a discretization of atmospheric flow on a latitude circle. We simulate the system for $500$ time units with the time step size $\Delta t=10^{-3}$ with the above parameters. With a sub-sampling time step size $\Delta t=0.01$, the first $300$ time units are used as training data, and the rest $200$ time units are used for testing. Part of the spatial-temporal simulation results of the true system including the time series, the PDF, and the ACF of the state variable $x_{2}$, are shown in Fig. 3.3. Since the state variables in this Lorenz 96 system are statistically homogeneous, i.e., each state variable possesses identical statistical properties, the results of other state variables closely resemble the ones of $x_{2}$ shown in Fig. 3.3. The PDF and ACF are calculated based on all the 300 time units of training data. From the Hovmoller diagram in Fig. 3.3(a), we can see the system is chaotic and has wave patterns. In this numerical test case, the Lorenz 96 system is partially observed with the observed state variables being odd indexed, i.e., $\mathbf{u}_{1}=[x_{1},x_{3},\cdots,x_{39}]^{\mathtt{T}}$, and the unobservable ones being even indexed, i.e., $\mathbf{u}_{2}=[x_{2},x_{4},\cdots,x_{40}]^{\mathtt{T}}$.
Figure 3.3: Simulation results and statistics of the Lorenz 96 system, with (a) the Hovmoller diagram of the spatio-temporal patterns, (b) time series of $x_{2}$, (c) the PDF of $x_{2}$, and (d) the ACF of $x_{2}$. Note that the PDF and ACF are estimated from a simulation longer than the one presented in panel (b). The behavior at other state variables is similar to $x_{2}$ since the system is statistically homogeneous.
@@
In addition to the statistical homogeneity, locality accounts for another feature of the Lorenz 96 system, i.e., the dynamics of a state variable only depend on its nearby state variables. In this work, locality and homogeneity are both incorporated into the models if possible. More specifically, we assume that $\mathop{}\!\mathrm{d}x_{i}/\mathop{}\!\mathrm{d}t$ only depends on the local state variable $x_{i}$ and two nearby state variables $[x_{i-2},x_{i-1},x_{i+1},x_{i+2}]^{\mathtt{T}}$. On the other hand, the homogeneity indicates that the model that approximates the dynamics $\mathop{}\!\mathrm{d}x_{i}/\mathop{}\!\mathrm{d}t$ based on its dependent state variables is consistent for every $i$. It is worth noting that the models with conditional Gaussian structure only allow incorporating the homogeneity of the model form for the observed state variables and the unobservable ones separately, while the training data can still inform the models that all state variables of the true system are statistically homogeneous.

### v2

@@ 3 Numerical Experiments
    The true model. This corresponds to the governing equation of the complex dynamical systems in Eq. (LABEL:eq:DS_u1u2). State predictions are made by averaging ensemble simulations from any given initial state, while DA results are obtained using the ensemble Kalman-Bucy filter (EnKBF) [51] based on data from observed states. In this work, the EnKBF is employed exclusively by this model for DA, while other models use the analytic DA formulae from Eq. (LABEL:eq:CGKN_CGF) to ensure a fair comparison. The test results from the true model serve as a benchmark for optimal performance, though they may not always be the best in practical applications.
@@
    Standard Koopman network (Standard KoopNet). This model applies the standard Koopman theory to the dynamics of the unobserved state $\mathbf{u}_{2}$, resulting in a linear governing equation for the latent state variables $\mathbf{v}$, rather than the nonlinear but conditionally linear structure of $\mathbf{v}$ found in the CGKN model. It is worth highlighting that this model still retains a strongly nonlinear structure in the dynamics of $\mathbf{u}_{1}$. The detailed form of the standard Koopman network is:
@@
    $\frac{\mathop{}\!\mathrm{d}\mathbf{v}}{\mathop{}\!\mathrm{d}t}=\mathbf{F}_{2}+\mathbf{G}_{2}\mathbf{v}+{\bm{\sigma}}_{2}\dot{\mathbf{W}}_{2},$
    where $\mathbf{F}_{2}\in\mathbb{R}^{d_{\mathbf{v}}}$ and $\mathbf{G}_{2}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ are two parameter matrices with other components staying the same as the CGKN model in Eq. (2.8). This standard application of the Koopman operator remains to impose a conditional Gaussian structure on the model, but the linear governing equation for $\mathbf{v}$ overlooks the contribution of the observed state $\mathbf{u}_{1}$. Numerical results show that, in comparison to the standard KoopNet model, the CGKN model’s conditional linear structure in the governing equation of $\mathbf{v}$ can further enhance its expressiveness. The comparison between the CGKN and the standard KoopNet model highlights the crucial role of the conditional linear, yet still nonlinear, structure in the governing equation of $\mathbf{v}$ in the former. The standard KoopNet can be viewed as a simplified version of CGKN where the latent dynamics do not explicitly depend on the observed states. If the true dynamical system possesses this characteristic, then the expressiveness of the standard KoopNet is sufficient for it to perform comparably to CGKN.
@@
    Conditional Gaussian regression (CG-Reg). This is a regression model from system identification that involves two steps: (i) identifying significant basis functions from a library of system dynamics using a causal-based metric called causation entropy [99, 100, 101, 102, 103, 104], and (ii) estimating the coefficients of these significant basis functions via the least squares method or maximum likelihood estimation. The library is specifically designed to ensure that the resulting regression model conforms to the structure of the conditional Gaussian nonlinear system in Eq. (2.8), allowing the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) to still be applied for DA [21]. In this work, this specially designed library is referred to as the CG-library, which requires that the unobserved state variables $\mathbf{u}_{2}$ appear linearly in all the basis functions. The detailed form of the CG-Reg model is as follows:
@@
    where $\mathbf{u}$ is the original system state variables including both the observed and unobserved ones. This DNN model does not follow the conditional Gaussian structure, and therefore, the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) cannot be applied for efficient DA. Numerical results demonstrate that the DNN model achieves state forecast performance comparable to that of the CGKN model; however, the additional key advantage of the CGKN model lies in its unique efficiency in DA.
For all numerical test cases, we assume that both the training and testing data include observed and unobserved state variables, although the inference of a trained CGKN does not require unobserved state variables. To evaluate the performance of models in both state prediction and DA across all examples, the normalized root mean squared error (NRMSE) is used:
$\textrm{NRMSE}:=\frac{1}{M}\sum_{i=1}^{M}\frac{\sqrt{\frac{1}{N}\sum_{n=1}^{N}\big{(}\mathbf{x}_{i}^{\star}(t_{n})-\mathbf{x}_{i}(t_{n})\big{)}^{2}}}{\textrm{std}(\mathbf{x}_{i}^{\star})},$ (3.4)
where $\mathbf{x}^{\star}(t_{n})\in\mathbb{R}^{M}$ is the true variables at the time step $t_{n}$, $\mathbf{x}(t_{n})$ is the approximated variables, $\mathbf{x}_{i}(t_{n})$ denotes the $i$-th variable, and $\textrm{std}(\cdot)$ corresponds to the standard deviation estimated by the samples from different time steps $t_{n}=1,...,N$. The NRMSE between true state variables and predictive ones is referred to as the forecast error, and the NRMSE between the true state variables and estimated state variables from DA is referred to as the DA error. Both forecast error and DA error are calculated using test data. The forecast error is calculated based on the same total forecast time steps in training. In contrast, the DA error is calculated over the whole time range of the test dataset.
@@
The simulation results of 1000 time units are generated and provide the data with a sub-sampling step size $\Delta t=0.01$ time unit, where the first 800 units are utilized for training and the remaining 200 units are employed for testing. Figure 3.1 shows part of the simulation of PSBSE in panel (a), and the probability density functions (PDFs) and auto-correlation functions (ACFs) in panels (b) and (c), respectively. The selection of parameters plays a crucial role in the system’s behavior: the coefficient of linear terms $\beta_{x}=0.2$ introduces instability, while $\beta_{y}=-0.3$ and $\beta_{y}=-0.5$ linearly dampens the system. Additionally, setting $\alpha=5$ contributes significant non-linearity to the system. Under the parameter settings, the true system demonstrates several complex characteristics, including intermittency, extreme events, and non-Gaussian statistics, as illustrated in Figure 3.1. Among the three states, $x$ is set as the observed state, while $y$ and $z$ are set as unobserved states in this test case. This setup of observed and unobserved states aligns with real-world applications since the state $x$ is the largest scale variable in the system. Additional test results of different setups of observed and unobserved states can be found in A.
@@
where the latent state variables $\mathbf{v}$ are transformed from unobserved state variables $[y,z]^{\mathtt{T}}$ via the encoder $\bm{\varphi}$. The dimension of $\mathbf{v}$ is set as 10 in this test case, and the sub-networks in the CGKN model are $\bm{\varphi}:\mathbb{R}^{2}\mapsto\mathbb{R}^{10}$, $\bm{\psi}:\mathbb{R}^{10}\mapsto\mathbb{R}^{2}$, and $\bm{\eta}:\mathbb{R}^{1}\mapsto\mathbb{R}^{121}$ with number of parameters $\bm{\theta_{\bm{\varphi}}}$, $\bm{\theta_{\bm{\psi}}}$, and $\bm{\theta_{\bm{\eta}}}$ be 8778, 8770, and 5785 respectively. The fully connected neural networks are used in CGKN, with the encoder $\bm{\varphi}$, decoder $\bm{\psi}$ and neural networks $\bm{\eta}$ each consisting of five layers. The ReLU activation function is applied to all networks in CGKN. The training settings of the CGKN include: state prediction time range $[t_{0},t_{N_{s}}]$ is $0.1$ time units, and the DA time range $[t_{0},t_{N_{l}}]$ is $50$ time units with the warm-up period set as $t_{N_{b}}=3$ time units. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=\lambda_{\mathbf{v}}=1/d_{\mathbf{u}}$, and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is set as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler.
Table 3.1: Test results of state forecast and data assimilation of all models for the example of projected stochastic Burgers–Sivashinsky equation. The errors are the normalized root mean squared error (NRMSE) between true values and approximated values.
Models Errors Forecast Error DA Error True Model 2.4566e-01 6.9466e-01 CGKN 2.8389e-01 7.2776e-01 Standard KoopNet 3.7600e-01 7.5069e-01 CG-Reg 2.7766e-01 1.2884e+00 DNN 2.4728e-01
@@
We investigate the performance of both state prediction and DA of all models on the test data and report the results in Table 3.1. The forecast error is calculated by 0.1 time units as the forecast time range and the DA Error is based on 200 time units as the DA time range. For the true model, the ensemble number is set to 100 for both the ensemble mean forecast and the EnKBF. The error in the true model primarily arises from intrinsic random noise in the system, along with sampling errors of the ensemble methods. Consequently, the test results of the true model serve as a benchmark for evaluating the performance of other models. Among all other models, the CGKN model provides the best overall performance in both state forecast and DA. Although the DNN model slightly outperforms the CGKN model for state forecast, its black-box nonlinear architecture demands more computationally expensive DA techniques, while the CGKN model can exploit the analytic formulae of Eq. (LABEL:eq:CGKN_CGF) for efficient DA. The forecast error from the CG-Reg model is slightly better than that of the CGKN model. The main reason is that, except for the nonlinear $yz$ term, all other terms in the true system of Eq. (3.5) can still be successfully identified by the CG-Reg model if all those terms have been included in the CG library. On the other hand, the CG-Reg model provides the worst performance of DA, indicating that the nonlinear term of $yz$ in the governing equation of $x$ plays an important role in the probabilistic relation between the observed state variable $x$ and the unobserved ones $y$ and $z$. Regarding the standard KoopNet model, the state forecast error is expected to be larger than that of the CGKN model, primarily due to the omission of the contribution from the observed state variable $x$ in the governing equations of the unobserved state variables.
@@
Figure 3.2 presents the DA results from the four models for estimating the unobserved states $[y,z]^{\mathtt{T}}$ based on the observed state $x$. To obtain the posterior mean estimates, the EnKBF is employed for the true model, while the analytic formulae in Eq. (LABEL:eq:CGKN_CGF) are used for the other three models. The uncertainty regions for the true model and the CG-Reg model are derived from the EnKBF and the covariance evolution formula in Eq. (LABEL:eq:CGKN_CGF), respectively. In contrast, the uncertainty regions for the CGKN and the standard Koopman network are obtained using the residual analysis method described in Section 2.2.2. The comparable performance of the posterior mean estimation to that of the true model underscores the effectiveness of the CGKN in DA. Furthermore, the estimated uncertainty region encompasses most of the results from the true model, even though it remains narrower than that produced by the true model with EnKBF. In this numerical test case, the DA results from the standard KoopNet model shown in Fig. 3.2(c) are comparable to those of the CGKN model, while the results from the CG-Reg model in Fig. 3.2(d) are noticeably inferior to those of the other models, particularly for the unobserved state variable $z$.
@@
Figure 3.2: DA results of the projected stochastic Burgers–Sivashinsky equation. EnKBF is used for the true model and analytic formulae are used for the other three models. The uncertainties are indicated by the grey-colored regions, which correspond to two estimated standard deviations from the posterior mean. The uncertainty area of the CGKN model and the standard KoopNet model are estimated based on the method of residual analysis.
@@ 3.2 Lorenz 96 system: A forty-dimensional chaotic system
where $I=40$, $F=8$ and $\sigma=0.5$. Under the specified parameter settings, the Lorenz 96 system exhibits chaotic and wave-like behaviors, allowing it to be regarded as a discretization of atmospheric flow on a latitude circle. Therefore, the current choice of parameters makes the system particularly useful for studying weather patterns and atmospheric dynamics in a controlled yet realistic setting. We simulate the system for $500$ time units with the time step size $\Delta t=10^{-3}$ with the above parameters. With a sub-sampling time step size $\Delta t=0.01$, the first $300$ time units are used as training data, and the rest $200$ time units are used for testing. Part of the spatial-temporal simulation results of the true system, including the time series, the PDF, and the ACF of the state variable $x_{2}$, are shown in Fig. 3.3. Since the state variables in this Lorenz 96 system are statistically homogeneous, i.e., each state variable possesses identical statistical properties, the results of other state variables closely resemble the ones of $x_{2}$ shown in Fig. 3.3. The PDF and ACF are calculated based on all the 300 time units of training data. In addition to the statistical homogeneity, locality accounts for another feature of the Lorenz 96 system, i.e., the dynamics of a state variable only depend on its nearby state variables. In this numerical test case, a classic DA setup is employed where every second state of the Lorenz 96 system was designated unobserved. More specifically, the system is partially observed with the observed state variables being odd indexed, i.e., $\mathbf{u}_{1}=[x_{1},x_{3},\cdots,x_{39}]^{\mathtt{T}}$, and the unobserved ones being even indexed, i.e., $\mathbf{u}_{2}=[x_{2},x_{4},\cdots,x_{40}]^{\mathtt{T}}$. From this setup, we demonstrate that the physics information, including homogeneous property and local spatial structure, can be compiled into CGKN, which forces the model to be consistent with prior conditions and reduces computational costs.
@@
In this work, locality and homogeneity are both incorporated into the models if possible. More specifically, we assume that $\mathop{}\!\mathrm{d}x_{i}/\mathop{}\!\mathrm{d}t$ only depends on the local state variable $x_{i}$ and two nearby state variables $[x_{i-2},x_{i-1},x_{i+1},x_{i+2}]^{\mathtt{T}}$. On the other hand, the homogeneity indicates that the model that approximates the dynamics $\mathop{}\!\mathrm{d}x_{i}/\mathop{}\!\mathrm{d}t$ based on its dependent state variables is consistent for every $i$. It is worth noting that the models with conditional Gaussian structure only allow incorporating the homogeneity of the model form for the observed state variables and the unobserved ones separately, while the training data can still inform the models that all state variables of the true system are statistically homogeneous.
Figure 3.3: Simulation results and statistics of the Lorenz 96 system, with (a) the Hovmoller diagram of the spatiotemporal patterns, (b) time series of $x_{2}$, (c) the PDF of $x_{2}$, and (d) the ACF of $x_{2}$. Note that the PDF and ACF are estimated from a simulation longer than the one presented in panel (b). The behavior of other state variables is similar to $x_{2}$ since the system is statistically homogeneous.